Por favor me ayuda a factorizar esta expresion: x^3-1
Por favor me ayuda a factorizar esta expresion: 2y^2+9-18
Por favor me ayuda a factorizar esta expresion: 8y^3+12y^2-18y-27

Answer by rothauserc(4718)

1)x^3 - 1 = (x -1)(x^2 +x +1)
Esta es la diferencia de dos cubos
2) 2y^2 +9 -18 = 2y^2 -18 +9 = (2y +6)(y -3) +9
3) 8y^3 +12y^2 -18y -27 = 4y^2(2y +3) -9(2y+3) = (4y^2 -9)(2y +3) =
(2y +3)(2y -3)(2y +3) = (2y +3)^2(2y -3)